# Computation of stresses in jammed packings modeled with Tresca friction

### v1

@@ Abstract
This paper is interested in the computation of stresses within jammed packings of rigid polygonal cells. The cells are considered to follow a Tresca friction law. First, a constrained minimization problem is introduced where the friction energy is minimized while enforcing the non-interpenetration of neighboring cells as inequality constraint. The corresponding dual maximization problem is then deduced and its solution provides normal stresses at the interface between cells. Finally, lowest order Raviart–Thomas finite elements are used to reconstruct a consistent stress field by solving local problems. Numerical results are presented to showcase the consistency and robustness of the proposed methodology.
@@ 1 Introduction
Granular materials inherently contain a lot of voids, which is an extra difficulty on its own. This paper is concerned with jammed packings that do not have such voids e.g. brick walls or vaults, see Figure 1.
@@
Let us denote by $\Omega\subset\mathbb{R}^{d}$, where $d=2,3$ a domain of interest (the jammed packing). We will consider that a packing is defined by a Voronoi mesh, denoted by $\mathcal{M}$, which perfectly matches the domain $\Omega$, see Figure 2.
Figure 2: Voronoi mesh $\mathcal{M}$ of $\Omega$.
Throughout this paper, we will refer to each element of the packing as a cell (of the mesh $\mathcal{M}$). In Figure 2, the blue dots represent the Voronoi centers of each cell. The interface between cells will then be the internal edges of the mesh $\mathcal{M}$.
@@
Since [5], many DEM have computed friction forces by penalizing interpenetration. The friction forces are then deduced linearly from the interpenetration. Plasticity can also be taken into account in [5]. More recently, [13] has proposed an approach, using DEM, that does not allow interpenetration of neighboring cells and does not use elastic parameters. Within appropriate loading regimes, we believe this approach better captures the underlying physics of jammed packings. The main difference with our work is that [13] models large displacements whereas we are interested in the stresses of jammed packings.
@@ 1.1 Tresca Friction
Let us illustrate the Tresca friction law in a very simple case. One can refer to the reference [3] for further detail. Let $\Omega\subset\mathbb{R}^{2}$ be an open bounded domain. We assume that the boundary of $\Omega$, written $\partial\Omega$, is Lipschitz and write as $\Gamma_{c}\subset\partial\Omega$ the contact surface. We write as $\bm{n}$ the outward unit normal to $\partial\Omega$ and as $\bm{t}=R(\frac{\pi}{2})\bm{n}$ the unit tangent to $\partial\Omega$, where $R(\frac{\pi}{2})$ is the 2d rotation matrix by an angle $\frac{\pi}{2}$. We write as $\bm{u}:\Omega\to\mathbb{R}^{2}$ the displacement field and $\sigma:\Omega\to\mathbb{R}^{2\times 2}_{\text{sym}}$ the stress tensor. We consider that the material in $\Omega$ is potentially in contact at $\Gamma_{c}$ with a fully rigid material (it does not deform). See figure 3 for a sketch of the setup.
@@ 2 Discrete Model
We consider an open bounded polyhedral domain $\Omega\subset\mathbb{R}^{2}$. For simplicity, the domain will be subjected to normal stress (or Neumann) boundary conditions but no volume loads. Include volume loads would be a straightforward extension.
@@ 2.1 Setup
Let $\mathcal{C}$ be a collection of Voronoi cells perfectly matching $\Omega$. Let us write as $\mathcal{E}$ the set of edges of the mesh. The edges are partitioned as $\mathcal{E}=\mathcal{E}^{\partial}\cup\mathcal{E}^{i}$, where $\mathcal{E}^{\partial}$ are the external edges and $\mathcal{E}^{i}$ are the internal edges. For a cell $c\in\mathcal{C}$, we write the position of its Voronoi center by $\bm{x}_{c}$, which also represents its initial position. Each cell $c\in\mathcal{C}$ is composed of a collection of edges written $\mathcal{E}_{c}$. The cells are considered to be rigid, and thus we take the current position $\bm{z}\in\mathbb{P}^{0}(\mathcal{C})^{2}=:V$ as piecewise constant in each cell. The initial position of the cells is also represented by a piecewise constant functions $\bm{x}\in V$. The displacement $\bm{u}_{c}$ of a cell $c\in\mathcal{C}$ is defined as $\bm{u}_{c}:=\bm{z}_{c}-\bm{x}_{c}$. The collection of cell displacements is then stored in $\bm{u}\in V$. Since the cells are considered as rigid bodies, a complete description of the kinematics of cell $c$ would require a translation of the centroid, $\bm{u}_{c}$ and a rotation $\theta_{c}$, about that point. But, since in the present context, we consider jammed packings, the cells cannot rotate and we therefore do not consider any rotational degree of freedom. As a consequence, only force balance is imposed on each cell, while moment balance is not enforced.
Let us consider two cells $c_{+}$ and $c_{-}$, such that $e\in\mathcal{E}_{c_{+}}\cap\mathcal{E}_{c_{-}}$, as in Figure 4.
$\bm{x}_{c_{+}}$$\bm{x}_{c_{-}}$$\bm{n}_{e}$$e$
Figure 4: Notation for edge quantities
For $e\in\mathcal{E}^{i}$, one defines $\bm{n}_{e}$ as the unit normal to $e$ pointing from $c_{-}$ towards $c_{+}$. The jump operator is defined for $\bm{u}$ at $e$ by
$\bm{u}]_{e}:=\bm{u}_{c_{+}}-\bm{u}_{c_{-}}.$
@@ 2.3 Primal problem
We consider that the cells interact through Tresca friction forces, and that there can be no inter-penetration between the cells. Let us first describe the energy $E_{e}$ between the two neighboring cells $c_{1}$ and $c_{2}$ of Figure 4. We write the energy of the Tresca friction law (compare to (3)) for $e$ as
@@ 2.4 Dual problem
Note that typical external loads $\bm{g}$ verifying the hypothesis of Proposition 1 contain compressive loads.
@@
$\max_{\bm{f}\in W^{2}}\mathtt{E}^{*}(\bm{f})$ (13)
Note that, because $\mathtt{E}$ has the physical dimension of an energy, then $\bm{f}$ is interpreted as a collection of forces associated to each internal edge, with the corresponding physical dimension. Let us recall the following classical result.
@@
Let $e\in\mathcal{E}^{i}$, we write $\bm{f}_{e}=(f^{n}_{e},f^{t}_{e})$. Assuming (13) has a feasible solution, the normal stress for a cell $c\in\mathcal{C}$ and an internal edge $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$ is thus defined as
@@
The principle of action-reaction is intrinsically taken into account here since for $c,c^{\prime}\in\mathcal{C}$ and $e\in\mathcal{E}_{c}\cap\mathcal{E}_{c^{\prime}}$, one has $\bm{\lambda}_{e,c}=-\bm{\lambda}_{e,c^{\prime}}$. That is due to the fact that one has, for instance, $\bm{n}_{c}=\bm{n}_{e}$ and $\bm{n}_{c^{\prime}}=-\bm{n}_{e}$.
@@ 3 Stress reconstruction
Solving (13) gives us, through (14), a normal stress in each internal edge $e\in\mathcal{E}^{i}$, which are completed by the normal stresses $\bm{g}$ imposed as Neumann boundary conditions on $\mathcal{E}^{\partial}$. Let us consider a cell $c\in\mathcal{C}$ as in Figure 5.
@@
We define as $H(\mathrm{div};c)^{2}:=\{\sigma\in L^{2}(c)^{2\times 2};\mathrm{div}(\sigma)\in L^{2}(c)^{2}\}$. The associated norm is defined for $\sigma\in H(\mathrm{div};c)$ by $\|\sigma\|_{H(\mathrm{div};c)}^{2}:=\|\sigma\|_{L^{2}(c)}^{2}+\|\mathrm{div}(\sigma)\|_{L^{2}(c)}^{2}$. The idea of the reconstruction is to compute in each cell $c\in\mathcal{C}$, a field $\sigma_{c}\in H(\mathrm{div};c)^{2}$, so that $\mathrm{div}(\sigma_{c})=\bm{0}$ in $c$, since no volume load is applied to $\Omega$. This is performed by solving local problems with finite elements, as described in the following. Note that, if volumes loads were applied to $\Omega$, one could modify the continuity equation in order to take the volumes loads into account.
@@ 3.1 Geometric setup
Let $c\in\mathcal{C}$. We write as $\bm{n}_{c}$ the outward unit normal to $\partial c$ and the barycentre of the internal edge $e\in\mathcal{E}^{i}$ as $\bm{x}_{e}$. Since all cells are Voronoi, they are convex. As a consequence, its boundary $\partial c$ is Lipschitz [8]. Since any cell $c\in\mathcal{C}$ is convex, it can be triangulated by choosing an edge $e\in\mathcal{E}_{c}$ and the Voronoi center $\bm{x}_{c}$, as sketched in Figure 5. Let us write as $\mathcal{M}_{c}$ this triangulation of the cell $c$.
@@ 3.2 Discrete problem
Note that Proposition 6 shows that from a given $\bm{\lambda}\in W^{2}$, solution of (13), one can compute a unique $\sigma\in\mathfrak{V}_{D}$. But since $\bm{\lambda}\in W^{2}$, solution of (13) is not unique, the reconstructed stress $\sigma$ is not unique either.
@@
In order to prove that (15) admits a unique solution, we want to use the BNB theorem for mixed systems, see [6, Theorem 2.34]. Let us first lift the Dirichlet boundary conditions. Using the surjectivity of the trace operator, see [7, Theorem 4.15], there exists $\mu\in H(\mathrm{div};c)^{2}$, such that for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{\partial}$, $\int_{e}\mu\bm{n}=|e|\bm{g}_{e}$ and for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$, $\int_{e}\mu\bm{n}=|e|\bm{\lambda}_{e,c}$. For $\tau\in\mathfrak{V}_{0}$, let us define the linear form
@@
Using [4, Theorem 2.3] and the fact that $\mathrm{curl}(\varphi)=\bm{0}$, one has
@@
We now consider a sequence $(\varphi_{n})_{n}\in C^{\infty}_{c}(c)^{2\times 2}$, such that $\mathrm{curl}(\varphi_{n})=\bm{0}$ and $\sigma\in Z$ such that $\|\varphi_{n}-\sigma\|_{H(\mathrm{div};c)}\to 0$, as $n\to+\infty$. Thus, as $n\to+\infty$, $\|\varphi_{n}-\sigma\|_{L^{2}(c)}\to 0$ and $\|\mathrm{div}(\varphi_{n}-\sigma)\|_{L^{2}(c)}$. One has
@@
In order to apply the BNB theorem for mixed formulation [6, Theorem 2.34], it remains to show the second condition of the theorem. Let $q\in Q$ and $\tau\in\mathfrak{V}_{0}$. One has
$b(\tau,q)=\sum_{T\in\mathcal{M}_{c}}\int_{T}\tau:j(q)$ $=\sum_{T\in\mathcal{M}_{c}}|T|\left<\tau\right>_{T}:j(q),$
$\leq\sum_{T\in\mathcal{M}_{c}}|T|\left|\left<\tau\right>_{T}\right|\sqrt{2}|q|,$
$\leq\sqrt{2}\left(\sum_{T\in\mathcal{M}_{c}}|T|\left|\left<\tau\right>_{T}\right|^{2}\right)^{\frac{1}{2}}\left(\sum_{T\in\mathcal{M}_{c}}|T||q|^{2}\right)^{\frac{1}{2}},$
$=\sqrt{2}\|\left<\tau\right>_{T}\chi_{T}\|_{L^{2}(c)}\|q\|_{L^{2}(c)}$
where $\left<\tau\right>_{T}$ is the average over $T$ and $\chi_{T}$ is the indicator function over $T$. Considering $\tau\in\mathfrak{V}_{0}\cap\mathbb{P}^{0}(\mathcal{M}_{c})^{2\times 2}$, one has $\|\tau\|_{H(\mathrm{div};c)}=\|\tau\|_{L^{2}(c)}$, and thus
$\frac{b(\tau,q)}{\|\tau\|_{H(\mathrm{div};c)}}=\sqrt{2}\|q\|_{L^{2}(c)}.$
The supremum over $\tau\in\mathfrak{V}_{0}$ is thus achieved and one has
$\inf_{q\in Q}\sup_{\tau\in\mathfrak{V}_{0}}\frac{b(\tau,q)}{\|\tau\|_{H(\mathrm{div};c)}\|q\|_{L^{2}(c)}}\geq\sqrt{2},$

### v2

@@ Abstract
This paper is interested in the computation of stresses within jammed packings of rigid polygonal cells. The cells are considered to follow a Tresca friction law. First, a constrained minimization problem is introduced where the friction energy is minimized while enforcing the non-interpenetration of neighboring cells as inequality constraints. The corresponding dual maximization problem is then deduced and its solutions provide normal stresses at the interface between cells. Finally, lowest order Raviart–Thomas finite elements are used to reconstruct a consistent stress field by solving local problems. Numerical results are presented to showcase the consistency and robustness of the proposed methodology.
@@ 1 Introduction
Granular materials inherently contain a lot of voids, which is an extra difficulty on its own. This paper is concerned with dense, jammed packings of polygonal cells that do not have such voids e.g. brick walls or vaults, see Figure 1.
@@
Let us denote by $\Omega\subset\mathbb{R}^{d}$, where $d=2,3$ a domain of interest (the jammed packing). We will consider that a packing is defined by a mesh of convex polygonal cells, denoted by $\mathcal{M}$, see Figure 2.
Figure 2: Polygonal mesh $\mathcal{M}$ of $\Omega$.
Throughout this paper, we will refer to each element of the packing as a cell (of the mesh $\mathcal{M}$). In Figure 2, the blue dots represent the appropriate center for the cell (centroid for generic convex cells, Voronoi center for Voronoi cells, etc.). The interfaces between cells will then be the internal edges of the mesh $\mathcal{M}$.
@@
Since [6], many DEM have computed friction forces by penalizing interpenetration. The friction forces are then deduced linearly from the interpenetration. Plasticity can also be taken into account in [6]. More recently, [15] has proposed an approach, using DEM, that does not allow interpenetration of neighboring cells and does not use elastic parameters. Within appropriate loading regimes, we believe this approach better captures the underlying physics of jammed packings. The main difference with our work is that [15] models large displacements whereas we are interested in the stresses within jammed packings.
@@ 1.1 Tresca Friction
Let us illustrate the Tresca friction law in a very simple case. One can refer to [4] for further detail. Let $\Omega\subset\mathbb{R}^{2}$ be an open bounded domain. We assume that the boundary of $\Omega$, written $\partial\Omega$, is Lipschitz and write as $\Gamma_{c}\subset\partial\Omega$ the contact surface. We write as $\bm{n}$ the outward unit normal to $\partial\Omega$ and as $\bm{t}=R(\frac{\pi}{2})\bm{n}$ the unit tangent to $\partial\Omega$, where $R(\frac{\pi}{2})$ is the 2d rotation matrix by an angle $\frac{\pi}{2}$. We write as $\bm{u}:\Omega\to\mathbb{R}^{2}$ the displacement field and $\sigma:\Omega\to\mathbb{R}^{2\times 2}_{\text{sym}}$ the stress tensor. We consider that the material in $\Omega$ is potentially in contact at $\Gamma_{c}$ with a fully rigid material (it does not deform). See figure 3 for a sketch of the setup.
@@ 2 Discrete Model
We consider an open bounded polygonal domain $\Omega\subset\mathbb{R}^{2}$. For simplicity, the domain will be subjected to normal stress (or Neumann) boundary conditions but no volume loads. Including volume loads would be a straightforward extension.
@@ 2.1 Setup
Let $\mathcal{M}$ be a decomposition of a domain $\Omega$ into a collection of convex polygonal cells. Let us write as $\mathcal{E}$ the set of edges of the mesh. The edges are partitioned as $\mathcal{E}=\mathcal{E}^{\partial}\cup\mathcal{E}^{i}$, where $\mathcal{E}^{\partial}$ are the external edges and $\mathcal{E}^{i}$ are the internal edges. For a cell $c\in\mathcal{M}$, we denote the position of its center by $\bm{x}_{c}$, which also represents its initial position. Each cell $c\in\mathcal{M}$ is composed of a collection of edges written $\mathcal{E}_{c}$. The cells are considered to be rigid, and thus we take the current position $\bm{z}\in\mathbb{P}^{0}(\mathcal{M})^{2}=:V$ as piecewise constant in each cell. The initial position of the cells is also represented by a piecewise constant functions $\bm{x}\in V$. The displacement $\bm{u}_{c}$ of a cell $c\in\mathcal{M}$ is defined as $\bm{u}_{c}:=\bm{z}_{c}-\bm{x}_{c}$. The collection of cell displacements is then stored in $\bm{u}\in V$. Since the cells are considered as rigid bodies, a complete description of the kinematics of a cell $c\in\mathcal{M}$ would require a translation of a distinguished point $\bm{u}_{c}$, and a rotation $\theta_{c}$, about that point.
However, in our present modeling, we choose to neglect the rotational degrees of freedom. This simplification is motivated physically by the fact that the rotation of cells in a jammed configuration requires significant dilatancy and, as a consequence, the resistance of compressed packings to rotation is often orders of magnitude higher than the resistance to sliding (friction). This motivates our modeling approach where we solve only for force balance (dual to the translation degrees of freedom) and do not enforce the moment balance equations (which are the dual to the rotational degrees of freedom).
Remark 1.
In the context of frictionless packings, Voronoi cells are typically used because they automatically satisfy the moment equations. Without friction, the interaction forces are purely normal to the edges. Since the normal bisectors of a Voronoi cell all intersect at the center the net moment is always zero. For frictional packings, however, there is no particular mechanical advantage to using Voronoi cells over general polygonal cells, since tangential forces are present and moment balance is not automatically satisfied. We focus on convex cells here because they are common in practice and offer flexibility in choosing the cell center $\bm{x}_{c}$.
2.2 Hypertstatic packings
One of the key features of jammed packings is their static indeterminacy (or hyperstaticity). Let us count the degrees of freedom and constraints for a generic planar packing. The number of force balance equations depends on whether we enforce torque balance. Let $n$ be the number of force variables per edge ($n=1$ for frictionless, $n=2$ for frictional). Let $m$ be the number of moment balance equations per cell ($m=0$ if ignored/automatic, $m=1$ if enforced). For a system with $F$ cells and $J$ internal edges, the number of degrees of freedom is $nJ$ and the number of constraints is (2+m)F3(2+m)F-3 (subtracting 3 global rigid body modes). We summarize the different modeling choices in Table 1.
Model $\bm{n}$ (vars/edge) $\bm{m}$ (moments/cell) DoF Constraints
Frictionless Voronoi 1 0 (automatic) $J$ 2F32F-3
Frictional (Full Equilibrium) 2 1 $2J$ 3F33F-3
Frictional (Simplified, this paper) 2 0 $2J$ 2F32F-3
Table 1: Counting argument for different granular models. $n$: force variables per edge. $m$: moment equations per cell. Our model ($n=2,m=0$) is highly hyperstatic.
The condition for isostaticity would be DoF = Constraints. In our simplified frictional model ($n=2,m=0$), the number of variables $2J$ far exceeds the constraints 2F32F-3, leading to a large space of self-equilibrated stresses, corresponding to the dimension of the null space of the equilibrium matrix, s=nJ(2F3)>0s=nJ-(2F-3)>0. This means there exists a nontrivial subspace consisting of ‘states of self-stress’ that satisfy force balance with zero external load. Our optimization approach (10), and its numerical implementation effectively selects one specific solution from this hyperstatic manifold, and this state is one among the minimizers of the energy functional (8).
@@ 2.3 Facet values
For $e\in\mathcal{E}^{i}$, one defines $\bm{n}_{e}$ as the unit normal to $e$ pointing from $c_{-}$ towards $c_{+}$, see Figure 4. The jump operator is defined for $\bm{u}$ at $e$ by
$\bm{u}]_{e}:=\bm{u}_{c_{+}}-\bm{u}_{c_{-}}.$
$\bm{x}_{c_{+}}$$\bm{x}_{c_{-}}$$\bm{n}_{e}$$e$
Figure 4: Notation for edge quantities
@@ 2.4 Primal problem
We consider that the cells interact through Tresca friction forces, and that there can be no inter-penetration between the cells. Let us first describe the energy $E_{e}$ between the two neighboring cells $c_{+}$ and $c_{-}$ of Figure 4. We write the energy of the Tresca friction law (compare to (3)) for $e$ as
@@ 2.5 Dual problem
Since the energy (8) does not contain any elastic forces, see (5), one needs a confinement term $\bm{g}$ strong enough to avoid cells slipping to infinity. Section 4 contains several examples of such confinement terms $\bm{g}$.
@@
$\max_{\bm{f}\in W^{2}}\mathtt{E}^{*}(\bm{f}).$ (13)
Note that, since $\mathtt{E}$ has the physical dimension of an energy, $\bm{f}$ is interpreted as a collection of forces associated to each internal edge, with the corresponding physical dimension. Let us recall the following classical result.
@@
Remark 6.
In practice, if the solver used to compute a minimizer of (13) does not converge, adding a stronger confining term to $\bm{g}$ is often enough to obtain a solution of (13).
Let $e\in\mathcal{E}^{i}$, we write $\bm{f}_{e}=(f^{n}_{e},f^{t}_{e})$. Assuming (13) has a feasible solution, the normal stress for a cell $c\in\mathcal{M}$ and an internal edge $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$ is thus defined as
@@
The principle of action-reaction is intrinsically taken into account here since for $c,c^{\prime}\in\mathcal{M}$ and $e\in\mathcal{E}_{c}\cap\mathcal{E}_{c^{\prime}}\neq\emptyset$, one has $\bm{\lambda}_{e,c}=-\bm{\lambda}_{e,c^{\prime}}$. That is due to the fact that one has, for instance, $\bm{n}_{c}=\bm{n}_{e}$ and $\bm{n}_{c^{\prime}}=-\bm{n}_{e}$.
@@ 3 Stress reconstruction
Solving (13) gives us, through (14), a normal stress in each internal edge $e\in\mathcal{E}^{i}$, which are completed by the normal stresses $\bm{g}$ imposed as Neumann boundary conditions on $\mathcal{E}^{\partial}$. Let us consider a cell $c\in\mathcal{M}$ as in Figure 5.
@@
We define as $H(\mathrm{div};c)^{2}:=\{\sigma\in L^{2}(c)^{2\times 2};\mathrm{div}(\sigma)\in L^{2}(c)^{2}\}$. The associated norm is defined for $\sigma\in H(\mathrm{div};c)$ by $\|\sigma\|_{H(\mathrm{div};c)}^{2}:=\|\sigma\|_{L^{2}(c)}^{2}+\|\mathrm{div}(\sigma)\|_{L^{2}(c)}^{2}$. The idea of the reconstruction is to compute in each cell $c\in\mathcal{M}$, a field $\sigma_{c}\in H(\mathrm{div};c)^{2}$, so that $\mathrm{div}(\sigma_{c})=\bm{0}$ in $c$, since no volume load is applied to $\Omega$. This is performed by solving local problems with finite elements, as described in the following. Note that, if volumes loads were applied to $\Omega$, one could modify the continuity equation in order to take the volumes loads into account.
@@ 3.1 Geometric setup
Let $c\in\mathcal{M}$. We write as $\bm{n}_{c}$ the outward unit normal to $\partial c$ and the barycentre of the internal edge $e\in\mathcal{E}^{i}$ as $\bm{x}_{e}$. Since all cells are convex, the boundary $\partial c$ of $c\in\mathcal{M}$ is Lipschitz [10]. Also, since any cell $c\in\mathcal{M}$ is convex, it can be triangulated by choosing an edge $e\in\mathcal{E}_{c}$ and the center $\bm{x}_{c}$, as sketched in Figure 5. Let us write as $\mathcal{M}_{c}$ this triangulation of the cell $c$.
@@ 3.2 Discrete problem
Note that Proposition 8 shows that from a given $\bm{\lambda}\in W^{2}$, solution of (13), one can compute a unique $\sigma\in\mathfrak{V}_{D}$. But since $\bm{\lambda}\in W^{2}$, solution of (13), is not unique, the reconstructed stress $\sigma$ is not unique either.
Remark 10.
The use of $\mathbb{RT}_{0}$ elements implies a subtle consistency requirement regarding rotations. In standard elasticity, the symmetry of the stress tensor $\sigma=\sigma^{T}$ (which we impose weakly via (15)) is equivalent to the local balance of angular momentum (torque balance). However, since our discrete model (Section 2) neglects rotational equilibrium equations, the discrete forces/fluxes $\bm{\lambda}$ might not exactly sum to zero torque on each cell.
@@
In order to prove that (15) admits a unique solution, we want to use the BNB theorem for mixed systems, see [8, Theorem 2.34]. Let us first lift the Dirichlet boundary conditions. Using the surjectivity of the trace operator, see [9, Theorem 4.15], there exists $\mu\in H(\mathrm{div};c)^{2}$, such that for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{\partial}$, $\int_{e}\mu\bm{n}=|e|\bm{g}_{e}$ and for all $e\in\mathcal{E}_{c}\cap\mathcal{E}^{i}$, $\int_{e}\mu\bm{n}=|e|\bm{\lambda}_{e,c}$. For $\tau\in\mathfrak{V}_{0}$, let us define the linear form
@@
Using [5, Theorem 2.3] and the fact that $\mathrm{curl}(\varphi)=\bm{0}$, one has
@@
We now consider a sequence $(\varphi_{n})_{n}\in C^{\infty}_{c}(c)^{2\times 2}$, such that $\mathrm{curl}(\varphi_{n})=\bm{0}$ and $\sigma\in Z$ such that $\|\varphi_{n}-\sigma\|_{H(\mathrm{div};c)}\to 0$, as $n\to+\infty$. Thus, as $n\to+\infty$, $\|\varphi_{n}-\sigma\|_{L^{2}(c)}\to 0$ and $\|\mathrm{div}(\varphi_{n}-\sigma)\|_{L^{2}(c)}\to 0$. One has
@@
In order to apply the BNB theorem for mixed formulation [8, Theorem 2.34], it remains to show the second condition of the theorem. Let $q\in Q$ and $\tau\in\mathfrak{V}_{0}$. One has
$b(\tau,q)=\sum_{T\in\mathcal{M}_{c}}\int_{T}\tau:j(q)$ $=\sum_{T\in\mathcal{M}_{c}}|T|\mathrm{skew}(\left<\tau\right>_{T}):j(q),$
$\leq\sum_{T\in\mathcal{M}_{c}}|T|\left|\mathrm{skew}(\left<\tau\right>_{T})\right|\sqrt{2}|q|,$
$\leq\sqrt{2}\left(\sum_{T\in\mathcal{M}_{c}}|T|\left|\mathrm{skew}(\left<\tau\right>_{T})\right|^{2}\right)^{\frac{1}{2}}\left(\sum_{T\in\mathcal{M}_{c}}|T||q|^{2}\right)^{\frac{1}{2}},$
$=\sqrt{2}\left(\sum_{T\in\mathcal{M}_{c}}\|\mathrm{skew}(\left<\tau\right>_{T})\|^{2}_{L^{2}(T)}\right)^{\frac{1}{2}}\|q\|_{L^{2}(c)}$
where $\left<\tau\right>_{T}$ is the average over $T$, $\mathrm{skew}$ is the skew-symmetric part of a matrix, and we used the Cauchy–Schwarz inequality twice. Note that since $\tau$ is affine in each $T\in\mathcal{M}_{c}$, one has $\left<\tau\right>_{T}=\tau(\bm{x}_{T})$, for all $T\in\mathcal{M}_{c}$.
Let us now consider $\tau\in\mathfrak{V}_{0}$ such that for all $T\in\mathcal{M}_{c}$, $\mathrm{skew}(\tau(\bm{x}_{T}))=j(q_{|T})$. Let us explain why this is always possible when considering the kind of meshes $\mathcal{M}_{c}$ as in Figure 5. Let $n\in\mathbb{N}$ be the number of cells in $\mathcal{M}_{c}$. We thus have $6n$ dofs for $\tau\in\mathfrak{V}$. For $\tau\in\mathfrak{V}_{0}$, the boundary conditions imposed on $\partial c$ consist in $2n$ equations. There are as many cells as internal edges, see Figure 5. The normal continuity conditions at internal edges consist in $2n$ equations. Finally imposing $\mathrm{skew}(\tau(\bm{x}_{T}))=j(q_{|T})$, for all $T\in\mathcal{M}_{c}$, consists in $n$ equations. To conclude, we have $5n$ equations for $6n$ dofs.
With such a $\tau\in\mathfrak{V}_{0}$, the two Cauchy–Schwarz inequalities above are actually equalities. One thus has
$b(\tau,q)=\sqrt{2}\left(\sum_{T\in\mathcal{M}_{c}}\|\mathrm{skew}(\left<\tau\right>_{T})\|^{2}_{L^{2}(T)}\right)^{\frac{1}{2}}\|q\|_{L^{2}(c)}.$
Also, one has
$\sum_{T\in\mathcal{M}_{c}}\|\mathrm{skew}(\left<\tau\right>_{T})\|^{2}_{L^{2}(T)}\leq\sum_{T\in\mathcal{M}_{c}}\|\left<\tau\right>_{T}\|^{2}_{L^{2}(T)}\leq\sum_{T\in\mathcal{M}_{c}}\|\tau\|^{2}_{L^{2}(T)}=\|\tau\|^{2}_{L^{2}(c)}\leq\|\tau\|^{2}_{H(\mathrm{div};c)}.$
Therefore,
$\frac{b(\tau,q)}{\|\tau\|_{H(\mathrm{div};c)}}\leq\frac{b(\tau,q)}{\left(\sum_{T\in\mathcal{M}_{c}}\|\mathrm{skew}(\left<\tau\right>_{T})\|^{2}_{L^{2}(T)}\right)^{\frac{1}{2}}}=\sqrt{2}\|q\|_{L^{2}(c)}.$
The supremum over $\varphi\in\mathfrak{V}_{0}$ is thus achieved and one has
$\inf_{q\in Q}\sup_{\varphi\in\mathfrak{V}_{0}}\frac{b(\tau,q)}{\|\varphi\|_{H(\mathrm{div};c)}\|q\|_{L^{2}(c)}}\geq\sqrt{2},$